OpenStudy (anonymous):

@jim_thompson5910 A rectangular park measures 30 meters by 40 meters. How long will a diagonal path from one corner of the park to the other corner measure? A. 10 m B. 26.5 m C. 50 m D. 70 m Can you walk me through this?

OpenStudy (anonymous):

\[h^2=30^2+40^2\] solve for \(h\)

OpenStudy (anonymous):

notice that this is the same as a 3 - 4 - 5 right triangle only with 30 - 40 - ??

OpenStudy (anonymous):

I don't want the answer!!! I need to know this!!!! @CGGURUMANJUNATH Next time wait until so0meonew can figure it out on their own not give them the answer and then explain.

OpenStudy (anonymous):

it is a right triangle, and you are being asked for the hypotenuse as before

OpenStudy (anonymous):



OpenStudy (anonymous):

H^2 = 30^2 + 40^2 30^2 = 30 x 30 = 900 40^2 = 40 x 40 = 1600 C? 50? @satellite73

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

notice it is the mother of all right triangles \(3:4:5\) disguised as \(30:40:50\)

OpenStudy (anonymous):

seen any good youtube videos lately?

OpenStudy (anonymous):

Why yes I have. :) I got my computer fixed. :) Turns out I just downloaded something wrong. I <3 Bttle Scars by Paradise Fears. I would LOVE to hear Christina Grimmie sing it though. ;p

OpenStudy (anonymous):

you should make one!

OpenStudy (anonymous):

Me, make a video of me singing? Hah! I can't sing. I'm to quiet and I don't talk much. I'm more of a writer. I write my words not say them. :)